OpenStudy (anonymous):

solve for t 0=(-9.8/2)t^2+9t+300 please show step by step process solve for t 0=(-9.8/2)t^2+9t+300 please show step by step process @Mathematics

OpenStudy (anonymous):

rearrange the equation into the standard form, 0=a t^2+bt+c 0=-4.9t^2+9t+300 where a=-4.9, b=9 and c=300 and use the formula to get two answers, one by t={-b+sqrt[b^2-4*a*c]}/2 and the second by t={-b-sqrt[b^2-4*a*c]}/2

OpenStudy (anonymous):

there was a small error it should be and use the formula to get two answers, one by t={-b+sqrt[b^2-4*a*c]}/(2*a) and the second by t={-b-sqrt[b^2-4*a*c]}/(2*a)